# NovelQA: Benchmarking Question Answering on Documents Exceeding 200K Tokens

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has introduced a new frontier in natural language processing, particularly in understanding and processing long-context information. However, the evaluation of these models’ long-context abilities remains a challenge due to the limitations of current benchmarks. To address this gap, we introduce NovelQA, a benchmark specifically designed to test the capabilities of LLMs with extended texts. Constructed from English novels, NovelQA offers a unique blend of complexity, length, and narrative coherence, making it an ideal tool for assessing deep textual understanding in LLMs. This paper presents the design and construction of NovelQA, highlighting its manual annotation, and diverse question types. Our evaluation of Long-context LLMs on NovelQA reveals significant insights into the models’ performance, particularly emphasizing the challenges they face with multi-hop reasoning, detail-oriented questions, and extremely long input with an average length more than 200,000 tokens. The results underscore the necessity for further advancements in LLMs to improve their long-context comprehension and computational literary studies.
@@ 1 Introduction
Recent years have seen a remarkable surge in the development of Large Language Models (LLMs) (OpenAI, 2023a; Touvron et al., 2023). Among these developments, long-context LLMs stand out for their ability to process and interpret extended pieces of text (Tworkowski et al., 2023; Team, 2023; Anthropic, 2023). This capability is essential for complex tasks that require a deep and nuanced understanding of lengthy documents, such as legal cases Xiao et al. (2021) or academic papers Groeneveld et al. (2024), where the key is to understand extended narratives Xu et al. (2023). Meanwhile, the ability to analyze extremely long documents and multiple documents simultaneously is increasingly crucial, supporting more informed decision-making in various fields Deng et al. (2023); Lin et al. (2023); Boiko et al. (2023).
@@
However, evaluating the long-context capabilities of models presents challenges, as existing benchmarks (Yang et al., 2018; Tay et al., 2021) no longer align with the advanced processing abilities of current LLMs (Anthropic, 2023; Team, 2023; OpenAI, 2023b). The disparity is further highlighted by the increasing context window size of LLMs, which now outpaces the average token lengths found in long-range datasets. This gap is evident, as the most advanced long-context LLMs are capable of processing over 250,000 tokens, a stark contrast to the longest average token length in current benchmarks, which is around 60,000 tokens. This mismatch underscores the need for updated evaluation methods that can accurately reflect the capabilities of current and future LLMs, as illustrated in Figure 1.
@@
To fill this gap, we introduce NovelQA, a benchmark crafted to specifically evaluate LLMs’ performance on texts with context windows exceeding 200,000 tokens. Unlike existing benchmarks (Shaham et al., 2023; An et al., 2023; Adams et al., 2024), NovelQA addresses the need for assessing extremely long-context understanding, offering a refined and comprehensive tool for advancing natural language processing capabilities. We construct NovelQA based on novels in English, which are also ideal for testing long-context modeling because they are long and complex, with plots that are closely linked from start to end. We select novels from various eras, genres, and formats to enhance diversity. The annotation process is performed by a group of expert annotators, all of whom are holding or pursuing a degree in English Literature and have a strong interest in and familiarity with the novels they annotate. Each question is paired with a ‘golden answer’ and corresponding textual evidences from the novels. And we categorize them by complexity and aspect for detailed analysis. Figure 2 presents two examples, while Table 6 details the distribution of question types. The dataset includes multi-hop, single-hop, and detail questions, which test the model’s abilities to retrieve and integrate scattered information, retrieve information and summarize, and precisely identify specific and subtle details, respectively. The above capabilities are more challenging in a particularly long context. For example, finding relevant information is a particularly big challenge.
@@
Figure 2: Illustrative Examples from NovelQA: This figure showcases two sample questions. For each question, models are evaluated under two distinct settings – multi-choice, where the task is to select the correct answer from four options, and Generative, where the model generates an answer.
@@
Using the dataset, we assess various long-context LLMs, including commercial models such as GPT-4-128K (OpenAI, 2023b) and Claude 2.1-200K (Anthropic, 2023), alongside open-source models like InternLM2-Chat (Team, 2023). The primary findings offer a significant insight: even the most advanced long-context LLMs face challenges in consistently extracting and processing accurate information from extended texts. For example, GPT-4, the top performer, achieves a 46.88% accuracy rate, whereas the open-source InternLM2-20b achieves 32.37% in a generative setting. In addition, some models are unable to answer properly when the context exceeds 100K tokens, even if their context windows are much longer (xverse, 2023; 01-ai, 2023). Moreover, operating LLMs on inputs exceeding 200,000 tokens presents technical challenges, particularly regarding memory requirements and associated costs. The difficulty of valid models is particularly apparent in answering multi-hop questions and queries that probe meanings, relationships, spans, and timelines, highlighting a significant gap in the models’ long-range comprehension. Moreover, our data shows a decline in performance for evidence situated beyond the 100,000-token mark, including information at the novel’s end, diverging from the anticipated lost-in-middle phenomenon (Liu et al., 2023). This shift suggests a distinct challenge faced by LLMs when processing texts exceeding 100K tokens in length. These results highlight challenges not only in memory optimization but also in the nuanced comprehension and integration of lengthy texts, indicating a substantial obstacle on the path to truly effective long-context LLMs.
@@
To the best of our knowledge, NovelQA is the first long-context QA benchmark featuring manually crafted questions, golden answers, and evidences, with contexts extending beyond 200,000 tokens.  [^1] We have released the demonstrations and input of the testset, and created a leaderboard. More details can be found in https://github.com/NovelQA/novelqa.github.io. And NovelQA is released under the Apache-2.0 License.
@@ 2 Related Work
Long-Range Benchmarks Evaluating the ability of Long-Context Pretrained Language Models has been a hot topic (Tay et al., 2021; Shaham et al., 2023; Pang et al., 2022; Wang et al., 2024a; Yu et al., 2024). When entering the era of Large Language Models, the context window length has been much longer than ever, and the decoder-only LLMs have been the mainstream of language models and they have faced specific problems, such as the Lost-in-middle issue (Liu et al., 2023). Thus, increasing benchmarks are created for evaluating long-context LLMs. Among those benchmarks, several are representatives. The Needle in A Haystack test (gkamradt et al., 2023) involves inserting an unrelated sentence into a long context and asking the model to retrieve it. However, this test does not represent a natural language question answering task that requires composite and complex capabilities for processing long contexts. ZeroSCROLLS (Shaham et al., 2023), LooGLE (Li et al., 2023a), Counting-Stars(Song et al., 2024), and LongICLBench(Li et al., 2024) emphasize the importance of understanding and aggregating information from long texts, presenting challenges that highlight areas for improvement in current models. L-Eval (An et al., 2023) introduces a suite of tasks with human-labeled query-response pairs to assess LLMs’ performance in processing long inputs effectively, using advanced metrics for a more accurate evaluation. Furthermore, benchmarks like LongBench (Bai et al., 2023), BAMBOO (Dong et al., 2023a), LongBench-Chat (Bai et al., 2024) offer a diverse set of tasks across languages and domains, from reasoning and coding to summarization and multilingual translation. These benchmarks are designed to rigorously test the ability of LLMs to manage extensive contexts, with LongBench-Chat specifically focusing on instruction-following capabilities in long-context interactions. Additionally, LongHealth (Adams et al., 2024) addresses the need for LLMs to interpret long clinical documents accurately, providing a specialized benchmark for evaluating models on medical texts. This focus on domain-specific challenges underscores the broader necessity for LLMs to not only handle long texts but to do so in a manner that is accurate and contextually relevant across various fields.
@@
NovelQA distinguishes from existing benchmarks in three points. Firstly, NovelQA features an average token length exceeding 200,000, far surpassing the tens of thousands typically found in other benchmarks. Secondly, while other datasets often rely on LLMs or existing datasets for content creation, the questions, golden answers, and evidences of NovelQA are entirely crafted through human effort. We present a detailed comparison between NovelQA and other benchmarks in Appendix Fig 4.
@@
We also discuss related Long-Context Language Modeling methods in Appendix Sec 4.
@@ 3.1 Dataset Description
Data Formulation Every novel ($N$) in the dataset corresponds to multiple pieces of annotated data ($d_{i}$). Each piece of data consists of the following domains, question ($Q_{i}$), answer ($A_{i}$), multi-choices ($a_{i,0}$, $a_{i,1}$, $a_{i,2}$, and $a_{i,3}$), gold label ($a_{i,gold}$), evidences ($s_{i,0}$, $s_{i,1}$, …), and type ($Complx_{i}$ and $Aspect_{i}$), among which the answer domain is a short answer to the question, while the multi-choices and gold label domain forms multi-choices, indicating that our dataset can serve to the evaluation of both the generative task and the multi-choice task. In the generative setting, a novel text $N$ and a question $Q_{i}$ are combined to send into the model each time, and the generated answer is compared with the answer $A_{i}$. In the multi-choice setting, the novel $N$, a question $Q_{i}$, and the four choices $a_{i,0}$ to $a_{i,3}$ are sent into the model, and the output is evaluated according to the gold label $a_{i,gold}$. Meanwhile, the evidences domain consists of either the original excerpts from the novel, or the reasoning steps written by the annotator.
Book Selection We source public domain novels from Project Gutenberg, and purchase e-books from internet if it is necessary. We aim to enhance diversity by selecting novels from various eras, genres, and formats. Recognizing that some newer, popular novels are still under copyright protection, NovelQA inevitably includes some copyrighted novels, comprising a mixture of 65 novels in public domain and 24 ones in copyright-protection.  [^2] We evaluate all annotated questions in this paper and conduct analysis. For the public access, we have released all constructed data. However, we only release public domain novels. Therefore, we offer two types of metrics in the evaluation system and leaderboard: one for evaluating QAs within public domain novels and another for evaluating QAs across all novels. All selected books exceed 50K words (approximately 67k tokens) and are in English. A visualization of the distribution of book token lengths is provided in Appendix Sec D.2.
@@
Question Distribution The annotated questions can be classified by the complexity of solving the question and the aspect that the question focuses on. By complexity, The data are categorized into three complexity levels, multi-hop (35.0%), single-hop (42.8%), and detail (22.2%). The order of complexity is as follows: multi-hop > detail > single-hop. By the aspect that each question focuses on, the data entails seven types. A detailed specification of each type is listed in Appendix D.3. We have a total of 2305 questions, of which 1640 are from 65 public domain novels, while the remaining 665 are from 24 copyrighted novels. According to the classification above, the distribution of the questions in our dataset is displayed in Appendix Table 6. We also list the ability tested by each kind of question in Appendix Table 8.
3.2 Data Collection and Annotation
@@
Procedure Overview The annotation process is performed by a group of expert annotators.The annotation procedure consists of two phases: (1) Template-based phase: The annotators can fill entities into 10+ templates (see Appendix Sec D.1) that we design to be related to multi-hop or detailed information. This phase entails half of the data, mainly contributing to the multi-hop ones. (2) Free-formed phrase: To ensure the diversity of question expression and align the questions to the natural distribution, our second half of the data is annotated without a template, namely, the annotators contribute to any difficult questions that they come up with freely.
@@
Annotator Recruitment Our annotators are predominantly English Language and Literature university students or those with a keen interest in English novels, recruited from local universities. These students are accustomed to closely reading English novels for coursework or as a hobby, and writing reports or papers on them. Annotators are allowed to select novels for annotation based on their familiarity, ensuring they had previously read and comprehensively understood the texts. Meanwhile, we make sure the selected books meet our standards of enough word count and well-developed narratives. We also ensure that the each selected novel is either annotated by only one individual, or consistent in version across annotators, despite minor variations among different editions. Each annotator contribute to a typically small number of 20-30 questions per novel. This approach avoids forcing annotators annotating questions on unfamiliar content.
@@
Template Design The first annotation phase relies on a question template, which requires the annotator to fill in the entities from the novel to form valid questions. To design templates, we carried out sufficient pre-tests on GPT-4 and Claude2.1 to analyze their possible weaknesses in long-input QA and novel knowledge memorization. Our pre-test shows that they usually fail to tackle information spanning over multiple chapters, as well as lack attention to details that have no contribution to the main theme. Meanwhile, we also refer to around fifteen books on novel and narration theories (e.g. Forster, 1927; Tobias, 2012; Schmidt, 2012; McKee, 2005) to ensure our template covers more aspects that a novel can discuss (e.g., character, setting, theme).
Quality Control The created data is manually double-checked by three authors of this work. The review procedure follows the criteria of minimizing factual errors, enhancing expression clarity, and maintaining challenges to LLMs. Consequently, only 79.4% of the collected data are preserved, resulting a final dataset of 2305 QA tuples. Meanwhile, we have also conducted the inter-annotator agreement (IAA) test on the multi-choice versioned QA annotation on 100 cases, shows a score of 94.6% in Cohen’s Kappa, indicating a high agreement among annotators. We present the details of the IAA test in the Appendix Sec D.4.
Distractions for Multi-choice Setting We use GPT-4 to generate three distracting options for each question and its golden answer and randomly permute the four answers. And we check those distractions and rewrite those with similar meaning with the golden answers manually when we double check the data.
3.3 Advantages
Our NovelQA dataset serves as a new benchmark for evaluating long-context understanding, distinguished by several key advantages. Firstly, it surpasses existing benchmarks in length, offering a rigorous test of a model’s ability to navigate and comprehend significantly longer texts. Secondly, the inclusion of clear evidences alongside questions ensures that evaluations are grounded in concrete textual support, enhancing the reliability of assessments. Furthermore, the dataset emphasizes questions that require attention to detailed information, challenging models to move beyond superficial impressions to extract specific, nuanced answers. Questions, golden answers, and evidences of the dataset are entirely manually annotated and carefully checked, ensuring high-quality, nuanced questions and answers that reflect complex human thought processes. To prevent against data leakage, we will not release golden answers for the test set, minimizing the risk of overfitting. These features, combined with the dataset’s comprehensive coverage of diverse narratives and meticulous construction, make NovelQA a valuable resource for advancing long-context understanding.

### v3

@@ Abstract
Recent advancements in Large Language Models (LLMs) have pushed the boundaries of natural language processing, especially in long-context understanding. However, the evaluation of these models’ long-context abilities remains a challenge due to the limitations of current benchmarks. To address this gap, we introduce NovelQA, a benchmark tailored for evaluating LLMs with complex, extended narratives. Constructed from English novels, NovelQA offers a unique blend of complexity, length, and narrative coherence, making it an ideal tool for assessing deep textual understanding in LLMs. This paper details the design and construction of NovelQA, focusing on its comprehensive manual annotation process and the variety of question types aimed at evaluating nuanced comprehension. Our evaluation of long-context LLMs on NovelQA reveals significant insights into their strengths and weaknesses. Notably, the models struggle with multi-hop reasoning, detail-oriented questions, and handling extremely long inputs, with average lengths exceeding 200,000 tokens. Results highlight the need for substantial advancements in LLMs to enhance their long-context comprehension and contribute effectively to computational literary analysis.
@@ 1 Introduction
Recent years have seen a remarkable surge in the development of Large Language Models (LLMs) (OpenAI, 2023b; Touvron et al., 2023). Among these developments, long-context LLMs stand out for their ability to process and interpret extended pieces of text (Tworkowski et al., 2023; Team, 2023; Anthropic, 2023). This capability is essential for complex tasks that require a deep and nuanced understanding of lengthy documents, such as legal cases (Xiao et al., 2021) or academic papers (Groeneveld et al., 2024), where the key is to understand extended narratives (Xu et al., 2023). In addition, the ability to analyze extremely long documents and multiple documents simultaneously is increasingly crucial, supporting more informed decision-making in various fields (Deng et al., 2023; Lin et al., 2023; Boiko et al., 2023).
@@
The evaluation of extremely long-context capabilities presents challenges, as existing benchmarks (Yang et al., 2018; Tay et al., 2021) no longer align with the advanced processing abilities of current LLMs (Anthropic, 2023; Team, 2023; OpenAI, 2023a). The increasing context window size of LLMs now outpaces the average token lengths found in long-range datasets. This gap is evident, as the most advanced long-context LLMs are capable of processing over 250,000 tokens, a stark contrast to the longest average token length in current benchmarks, which is around 60,000 tokens. This mismatch underscores the need for updated evaluation methods that can accurately reflect the capabilities of current and future LLMs, as illustrated in Figure 1.
@@
To fill this gap, we introduce NovelQA, a benchmark crafted to specifically evaluate LLMs’ performance on texts with averaged context windows exceeding 200,000 tokens. Unlike existing benchmarks (Shaham et al., 2023; An et al., 2023; Adams et al., 2024), NovelQA addresses the need for assessing extremely long-context understanding, offering a refined and comprehensive tool for advancing natural language processing capabilities. We construct NovelQA based on novels in English, which are also ideal for testing long-context modeling because they are long and complex, with plots that are closely linked from start to end. We select novels from various eras, genres, and formats to enhance diversity. The annotation process is performed by a group of expert annotators, all of whom are holding or pursuing a degree in English Literature and have a strong interest in and familiarity with the novels they annotate. Each question is paired with a ‘golden answer’ and corresponding textual evidences from the novels. And we categorize them by complexity and aspect for detailed analysis. Figure 2 presents two examples, while Table 1 details the distribution of question types. The dataset includes multi-hop, single-hop, and detail questions, which test the model’s abilities to retrieve and integrate scattered information, retrieve information and summarize, and precisely identify specific and subtle details, respectively. The above capabilities are more challenging in a particularly long context. For example, finding relevant information is a particularly big challenge.
@@
Figure 2: Illustrative examples from NovelQA: This figure showcases two sample questions. For each question, models are evaluated under two distinct settings – multichoice, where the task is to select the correct answer from four options, and Generative, where the model generates an answer.
@@
We assess various long-context LLMs using NovelQA, including commercial models such as GPT-4-128K (OpenAI, 2023a) and Claude-2.1-200K, Claude-3-200K, Claude-3.5-200K(Anthropic, 2023), alongside open-source models like InternLM2-Chat (Team, 2023) and Llama-3.1 (Touvron et al., 2023). Results show that even the most advanced long-context LLMs face challenges in consistently extracting and processing accurate information from extended texts. For example, Claude-3.5, the top performer, achieves a 62.30% accuracy rate, whereas the open-source Llama-3.1 achieves 51.50% in a generative setting. In addition, some models are unable to answer properly when the context exceeds 100K tokens, even if their context windows are much longer (xverse, 2023; 01-ai, 2023). Technically, operating LLMs on inputs exceeding 200,000 tokens presents challenges, particularly regarding memory requirements and associated costs. The difficulty of valid models is particularly apparent in answering multi-hop questions and queries that probe meanings, relationships, spans, and timelines, highlighting a significant gap in the models’ long-range comprehension. Moreover, our data shows a decline in performance for evidence situated beyond the 100,000-token mark, including information at the novel’s end, diverging from the anticipated lost-in-middle phenomenon (Liu et al., 2023). This shift suggests a distinct challenge faced by LLMs when processing texts exceeding 100K tokens in length. These results highlight challenges not only in memory optimization but also in the nuanced comprehension and integration of lengthy texts, indicating a substantial obstacle on the path to truly effective long-context LLMs.
@@
To the best of our knowledge, NovelQA is the first long-context QA benchmark featuring manually crafted questions, golden answers, and evidences, with contexts extending beyond 200,000 tokens.  [^1] We have released the demonstrations and input of NovelQA, and created a leaderboard. More details can be found in https://novelqa.github.io/. And NovelQA is released under the Apache-2.0 License. For the public access, we have released all constructed data on Huggingface https://huggingface.co/datasets/NovelQA/NovelQA and an evaluation system on Codabench https://www.codabench.org/competitions/2727/. However, we only release public domain novels. Therefore, we offer two types of metrics in the evaluation system and leaderboard: one for evaluating QAs within public domain novels and another for evaluating QAs across all novels.
@@ 2 Related Work
Long-Range Benchmarks Evaluating the ability of long-context Large Language Models has been a hot topic (Kočiský et al., 2018; Tay et al., 2021; Shaham et al., 2023; Pang et al., 2022; Wang et al., 2024b; Yu et al., 2024). When entering the era of Large Language Models, the context window length has been much longer than ever, and the decoder-only LLMs have been the mainstream of language models and they have faced specific problems, such as the Lost-in-middle issue (Liu et al., 2023). Thus, increasing benchmarks are created for evaluating long-context LLMs. Among those benchmarks, several are representatives. The Needle in A Haystack test (gkamradt et al., 2023) involves inserting an unrelated sentence into a long context and asking the model to retrieve it. However, this test does not represent a natural language question-answering task that requires composite and complex capabilities for processing long contexts. NarrativeQA (Kočiský et al., 2018) asks annotators from MTurk to read novel summaries to write questions. NarrativeXL (Moskvichev & Mai, 2023) utilizes GPT-3.5 to summarize approximately 150 scenes per book from a collection of 1,500 books, and automatically annotates around one million questions. ZeroSCROLLS (Shaham et al., 2023), LooGLE (Li et al., 2023a), Counting-Stars(Song et al., 2024), and LongICLBench (Li et al., 2024) emphasize the importance of understanding and aggregating information from long texts, presenting challenges that highlight areas for improvement in current models. L-Eval (An et al., 2023) introduces a suite of tasks with human-labeled query-response pairs to assess LLMs’ performance in processing long inputs effectively, using advanced metrics for a more accurate evaluation. Furthermore, benchmarks like LongBench (Bai et al., 2023), BAMBOO (Dong et al., 2023a), and LongBench-Chat (Bai et al., 2024) offer a diverse set of tasks across languages and domains, from reasoning and coding to summarization and multilingual translation. These benchmarks are designed to rigorously test the ability of LLMs to manage extensive contexts, with LongBench-Chat specifically focusing on instruction-following capabilities in long-context interactions. Additionally, LongHealth (Adams et al., 2024) addresses the need for LLMs to interpret long clinical documents accurately, providing a specialized benchmark for evaluating models on medical texts. This focus on domain-specific challenges underscores the broader necessity for LLMs to not only handle long texts but to do so in a manner that is accurate and contextually relevant across various fields.
@@
NovelQA distinguishes from existing benchmarks in three points. Firstly, NovelQA features an average token length exceeding 200,000, far surpassing the tens of thousands typically found in other benchmarks. Secondly, while other datasets often rely on AI-generation or existing datasets for content creation, the questions, golden answers, and evidences of NovelQA are entirely crafted through human effort of experts. We present a detailed comparison between NovelQA and other benchmarks in Appendix Fig 5.
@@
We also discuss related Long-Context Language Modeling methods in Appendix Sec C.1.
@@ 3.1 Dataset Overview
Data Formulation Every novel ($N$) in the dataset corresponds to multiple pieces of annotated data ($d_{i}$). Each piece of data consists of the following domains, question ($Q_{i}$), answer ($A_{i}$), multichoices ($a_{i,0}$, $a_{i,1}$, $a_{i,2}$, and $a_{i,3}$), gold label ($a_{i,gold}$), evidences ($s_{i,0}$, $s_{i,1}$, …), and type ($Complx_{i}$ and $Aspect_{i}$). NovelQA can serve for either a generative or a multichoice task. In the generative setting, a novel text $N$ and a question $Q_{i}$ are combined to send into the model each time, and the generated answer is compared with the answer $A_{i}$. In the multichoice setting, the novel $N$, a question $Q_{i}$, and the four choices $a_{i,0}$ to $a_{i,3}$ are sent into the model, and the output is evaluated according to the gold label $a_{i,gold}$. The evidences domain consists of either the original excerpts from the novel or the reasoning steps written by the annotator.
Book Selection The books in NovelQA contain 65 free public-domain books from the Project Gutenberg [^2] https://www.gutenberg.org/ and 24 copyright-protected books purchased from the Internet. The selection of the books follows a criterion to ensure the diversity in the eras, styles, and themes, thus we decided to include a portion of copyrighted novels. All selected books are in English and exceed 50K words (approximately 67k tokens). The distribution of the token count is illustrated in Figure 3.
Figure 3: Token Count Distribution in NovelQA, including Copyrighted (left) and Public Domain Only (right). The token counts of both the novel and the questions are counted. The tokenization procedure is accomplished through gpt-3.5-turbo-16k tokenizer.
Table 1: Distribution of Question Types in NovelQA: This table provides a breakdown of questions across different complexity categories (Multi-hop, Single-hop, Detail) and aspect categories (Times, Meaning, Span, Setting, Relation, Character, Plot).
Multi-hop Single-hop Detail Sum
Times 463 0 0 463
Meaning 34 126 206 366
Span 34 0 0 34
Setting 24 177 63 264
Relation 119 14 32 165
Character 69 255 98 422
Plot 64 414 113 591
Sum 807 986 512 2305
@@
Question Distribution The annotated questions can be classified by the complexity of solving the question and the aspect that the question focuses on. By complexity, the data are categorized into three complexity levels, multi-hop (35.0%), single-hop (42.8%), and detail (22.2%). The order of complexity is as follows: multi-hop > detail > single-hop. By the aspect that each question focuses on, the data entails seven types. A detailed specification of each type is listed in Appendix C.2.2. We have a total of 2305 questions, of which 1640 are from 65 public domain novels, while the remaining 665 are from 24 copyrighted novels. According to the classification above, the distribution of the questions in our dataset is displayed in Table 1. We also list the ability tested by each kind of question in Appendix Table 10.
Advantages Our NovelQA dataset serves as a new benchmark for evaluating long-context understanding, distinguished by several key advantages. Firstly, it surpasses existing benchmarks in length, offering a rigorous test of a model’s ability to navigate and comprehend significantly longer texts. Secondly, the inclusion of clear evidences alongside questions ensures that evaluations are grounded in concrete textual support, enhancing the reliability of assessments. Furthermore, the dataset emphasizes questions that require attention to detailed information, challenging models to move beyond superficial impressions to extract specific, nuanced answers. Questions, golden answers, and evidences of the dataset are entirely manually annotated and carefully checked, ensuring high-quality, nuanced questions and answers that reflect complex human thought processes. To prevent against data leakage, we will not release golden answers for the test set, minimizing the risk of overfitting. These features, combined with the dataset’s comprehensive coverage of diverse narratives and meticulous construction, make NovelQA a valuable resource for advancing long-context understanding.
3.2 Data Annotation
@@
Procedure Overview The annotation process is performed by a group of expert annotators. The annotation procedure consists of two phases: (1) Template-based phase: The annotators can fill entities into 10+ templates (see Appendix Sec C.2.1) that we design to be related to multi-hop or detailed information. This phase entails half of the data, mainly contributing to the multi-hop ones. (2) Free-form phrase: To ensure the diversity of question expression and align the questions to the natural distribution, the second half of our data is annotated without a template, namely, the annotators contribute to any difficult questions that they come up with freely.
@@
Annotator Recruitment Our annotators are predominantly English Language and Literature university students or those with a keen interest in English novels, recruited from local universities. These students are accustomed to closely reading English novels for coursework or as a hobby, and writing reports or papers on them. Before annotation, each annotator was instructed to read the annotation instructions to understand the requirements and sign to agree to participate. Annotators are allowed to select novels for annotation based on their familiarity, ensuring they have previously read and comprehensively understood the texts. Meanwhile, we make sure the selected books meet our standards of enough word count and well-developed narratives. We also ensure that each selected novel is either annotated by only one individual, or consistent in version across annotators, despite minor variations among different editions. Each annotator contributes to a typically small number of 20-30 questions per novel. This approach avoids forcing annotators to annotate questions on unfamiliar content.
Annotation Guideline While creating the QA tuples, our annotators are instructed to follow several principles below: (1) The annotators are either senior English language and literature college students, or students with high English test scores. (2) The annotators are required to read through the GPT-4 responses on example questions. (3) The annotators should choose novels above 50K tokens and read through the books they choose before annotation. (4) Evidence of each answer should be provided for validation purposes. Evidence should be as sufficient as possible.
Template Design The first annotation phase relies on a question template, which requires the annotator to fill in the entities from the novel to form valid questions. To design templates, we carried out sufficient pre-tests on several LLMs to analyze their possible weaknesses in long-input QA and novel knowledge memorization. Our pre-test shows that they usually fail to tackle information spanning over multiple chapters, as well as lack attention to details that have no contribution to the main theme. We also refer to around fifteen books on novel and narration theories (e.g. Forster, 1927; Tobias, 2012; Schmidt, 2012; McKee, 2005) to ensure our template covers more aspects that a novel can discuss (e.g., character, setting, theme). Templates are ensured to test on facts (e.g., events, entities, numbers) that can be traced back to specific evidences from the books, instead of on any subjective feelings or analysis of the readers.
@@
Quality Control The created data is manually double-checked by three authors of this work. The review procedure follows the criteria of minimizing factual errors, enhancing expression clarity, and maintaining challenges to LLMs. Besides, we ensured that all questions are based on factual descriptions and eliminated any subjective ones. Consequently, only 79.4% of the collected data are preserved, resulting in a final dataset of 2305 QA tuples. Meanwhile, we have also conducted the inter-annotator agreement (IAA) test, focusing on evaluating the quality of annotated question-answer pairs. Annotators are required to choose books they are familiar with but have not annotated themselves to answer questions on. As annotators are mostly from the same local university and share similar courses and projects, many can find books they have read in common. We select books read by at least two annotators and have the other reader answer multichoice questions. As the respondents are quite familiar with the target novels and have a strong academic background in English Literature, it takes them around 2-3 hours or less to complete each novel. The IAA test shows a score of 94.6% in Cohen’s Kappa, indicating a high agreement among annotators.
Distractions for Multichoice Setting We use GPT-4 to generate three distracting options for each question and its golden answer, and randomly permute the four answers. [^3] Our pilot study compared distractors generated by both GPT-4 and Claude 2.1. Interestingly, GPT-4 generated slightly more challenging distractors - both models scored approximately 0.5% lower on GPT-4’s distractors compared to Claude 2.1’s. This indicates no advantageous bias for GPT-4. We manually checked those distractions and rewrote those with similar meaning with the golden answers when we double check the data.